How do you solve this equation? Thank you so much.
7+3x/4 = -x/8

Answer by stanbon(75887)

7+3x/4 = -x/8
---
Note: It looks like the "4" is the denominator of "3x".
If so, multiply thru by 8 to get:
56 + 6x = -x
---
7x = -56
x = -8
=============
Cheers,
Stan H.